ninjabeeler:

A cylinder with a radius of 5 feet and height of 2 feet. V = B h. V = (5) (pi) (2) squared. V = 5 (pi) (4). V = 20 pi feet cubed. Amul is trying to find the volume of the cylinder. What was Amul’s error? Amul should have squared pi. Amul should not have multiplied by pi. Amul should have squared the radius, not the height. Amul did not make an error.

Florisalreadytaken:

the formula to find the volume of a cylinder is \[ V=B \times h ​\] so: \[ V= r^2 \times h \times\pi \] what Amul did is that she squared the height of it, not the radius

ninjabeeler:

gomawo thank u sry i speak korean